Please help. I looked over a bunch of your solutions already, but they are not the same questions.
Jack has part of $4500 invested at 4 percent and the rest at 6 percent. If his annual return on each investment is the same, then what is the average rate of interest, to the nearest tenth, that he will realize on $4500?
Thanks

Answer by jorel1380(3719)

.04n=.06(4500-n)
4n=6(4500-n)
4n=27000-6n
10n=27000
n=2700
Jack has 2700 invested at 4%, and 1800 at 6%.
2700x.04=108
1800x.06=108
216/4500=4.8%..